### v3

# ARCON: Advancing Auto-Regressive Continuation for Driving Videos

Abstract
Recent advancements in auto-regressive large language models (LLMs) have led to their application in video generation. This paper explores the use of Large Vision Models (LVMs) for video continuation, a task essential for building world models and predicting future frames. We introduce ARCON, a scheme that alternates between generating semantic and RGB tokens, allowing the LVM to explicitly learn high-level structural video information. We find high consistency in the RGB images and semantic maps generated without special design. Moreover, we employ an optical flow-based texture stitching method to enhance visual quality. Experiments in autonomous driving scenarios show that our model can consistently generate long videos.
1 Introduction
Figure 1: Auto-regressively generated minute-level video using ARCON. We show a sample video clip from the BDD100K dataset. We auto-regressively generate $45$ frames given the first $3$ frames at $0.6$Hz. The ego-car moves forward in a short period and changes lanes to the right in preparation for a right turn. After the right turn, it continues to move forward. This example demonstrates that our model can generate reasonable first-view driving videos and can generate a completely new scene after the turn.
Recently, we have witnessed the remarkable capability of auto-regressive large language models (LLMs) in generating high-quality text [63, 48, 30, 6, 45]. Many researchers have also attempted to convert data from other modalities into discrete tokens, aiming to leverage the success of LLMs [19, 1]. In the field of image generation, there is a significant amount of work developing along the paradigm of “next-token prediction”, such as VQVAE [62, 50], VQGAN [14], and MAGVIT [84]. They employ image tokenizers to transform continuous images into discrete tokens, and utilize auto-regressive models to generate image tokens. Recently, works such as MAGVIT-v2 [85], VAR [58], and LlamaGen [52] have further explored the upper limits of this research direction, achieving image generation results comparable to those of diffusion models [51].
In video-related research, video continuation or prediction [44, 42] is a task that is highly relevant to these studies. The learning objective of this task is considered to be key to building a world model [21, 22]. Currently, the most popular video generation works [40, 75, 77] are mainly based on the diffusion model. GameNGen [61] has achieved exceptional world modeling capability and stunning visual effects of predicted frames. Before the wave of LLMs, methods based on the auto-regressive models have shown potential in long-term motion prediction [74, 17]. However, efforts are still required to better adapt them to complex real-world scenarios.
Many researchers believe that LLMs have advantages in certain aspects, such as diverse generation, multi-modal fusion, and scalable model capability. Researchers aim to further harness the capabilities of the next-token prediction paradigm and explore scaling up to even larger models. Sequential Large Vision Model (LVM) [2] demonstrates that auto-regressive learning solely on tokenized image and video frame sequences can yield reasonable scaling performance and some non-trivial next-frame inference results. WorldGPT [18] acquires an understanding of world dynamics by analyzing millions of videos across various domains. Some cutting-edge works, such as VideoPoet [32] and GAIA-1 [25], have already achieved impressive results in token-based video generation.
These works inspire us to explore the potential of LVMs further in video continuation. The significant computational power and engineering investment required make exploration and ablation experiments challenging for LVM-based visual generation. There are currently many open problems that have not been fully researched:
  1. 1.
    How do we select a suitable tokenizer setting for videos, whose encoding tokens can be effectively learned by an auto-regressive model?
  2. 2.
    How to avoid degeneration into stationary results in long-term video generation?
  3. 3.
    How to improve the visual quality of generated results?
In this paper, we build a large auto-regressive transformer with up to $20$B parameters trained on large-scale tokenized video frames. We propose an ARCON scheme to utilize additional semantic tokens to help the model explicitly consider and learn about the structural information of the video. Thereby enhancing the temporal consistency and physical reasonableness of the video when auto-regressively generating very long videos. Fig. 1 shows a minute-long auto-regressive generated first-view driving video. Furthermore, we demonstrate that we can borrow some ideas from low-level vision methods [36, 90] to significantly improve the quality of the generated frames. We paste textures from the input high-resolution frames onto the low-resolution generated results through a cross-frame flow-based model. This method yields results that exceed expectations and is more convenient than improving the encoder of the tokenizer. We explore the video continuation task in autonomous driving scenarios and show impressive quantitative and qualitative experimental results.
Our work includes these main contributions:
  1. 1.
    We establish a video continuation model based on a visual tokenizer and LVM architecture that has the potential to achieve emergent capabilities.
  2. 2.
    We use semantic tokens and show they improve the model’s ability to create longer videos with better temporal consistency. The generated semantic maps have a good correspondence with the RGB images.
  3. 3.
    Our extensive experiments demonstrate that our model can produce long videos of diverse autonomous driving scenarios.
Figure 2: The structure of our ARCON model. left: We use Uniformer [33] to estimate the semantic maps. RGB images and semantic maps are encoded into discrete tokens using the same tokenizer [85]. right: We use an auto-regressive model to alternately predict RGB tokens and semantic tokens. During image decoding, the original frame can provide texture guidance for the generated results.
2 Related Work
Visual pre-training.
Supervised data is often expensive and limited in quantity, leading many researchers to focus on visual pre-training [23, 3]. This approach harnesses large models and unlabeled data to mirror the emergent abilities of LLMs, enabling a generalized model to tackle various visual tasks. ImageGPT [11] pioneers using next-pixel prediction to train an auto-regressive model for image generation, but struggled with high-resolution images due to direct pixel distribution learning. ViT [12] transforms visual tasks into sequence modeling by dividing images into patches and extracting embeddings, showcasing the transformer’s potential for computer vision, albeit limited to image classification. Subsequent works [4, 76, 23] adopt similar patch-based approaches, employing masked image modeling (MIM) for pre-training. They utilize visual tokenizers based on VQ-VAE [62], learned via auto-encoding, to convert image patches into tokens. During pre-training, masked patch tokens are predicted to reconstruct the original image. The Emu series [53, 54, 68] introduce a cross-modal tokenizer to uniformly encode text, images, and videos into tokens for large multi-modal model pre-training.
Visual token-based generation.
The visual generation task is currently a hot research area due to the impressive capabilities demonstrated by LLMs and the powerful generative abilities of diffusion-based models. However, few studies have focused on using discrete visual tokens for image and video generation, as diffusion models are believed to produce richer details. In contrast, auto-regressive models can result in image blurring due to information loss when quantizing continuous image features into discrete values via a visual tokenizer [55]. The core concept of these methods involves transforming images and videos from continuous pixel space to a discrete token space, enabling auto-regressive transformer models to handle visual sequence modeling. Notable works leveraging auto-regressive transformers for text-to-image generation include DALL-E [49], which employs a discrete variational autoencoder (dVAE) as an image tokenizer and generates image token sequences from text before decoding them back into images. To tackle the issue of low-quality images generated by auto-regressive models, HART [55] utilizes a diffusion model to explicitly recover detailed information lost during quantization, enhancing the quality of reconstructed images. Building on the diffusion transformer [46], Sora [5] achieves high-fidelity, long-sequence video generation, producing videos up to the minute level for the first time. Additionally, MAGVIT [84, 85] has shown that high-quality images and videos can be generated within a discrete token space, advancing the development of video generation based on discrete tokens [78, 64, 20, 53, 54, 68, 35, 2, 32, 72, 47]. These advancements highlight the potential of discrete visual tokens in vision generation tasks.
In-context visual learning.
The advantage of the video generation framework built based on a visual tokenizer and auto-regressive LLM is that it can convert most visual modalities (e.g., RGB image, segmentation map, depth map, optical flow, sketch, HED map) to the same token space, or even project modalities of different domains such as image, video, audio, text, etc., to the same embedding space for common modeling [19, 91], so that LLM can be used to translate between multiple modalities or use auxiliary modalities to control the generation of RGB videos [65, 66, 56, 43, 1, 69, 71, 82]. In our paper, we explore the interleaving generation of semantic tokens and RGB tokens to demonstrate that tokens containing highly structured information can aid language models in better understanding low-level RGB tokens therefore performing the video continuation task better.
Discrete visual tokenizer.
Discrete visual tokenizers [62, 13, 89, 10, 43, 86, 84, 85] have attracted increasing attention due to their strong linkage with LLMs and visual token-based generation. With a fixed-length vocabulary, which reduces the long-term attention complexity, discrete tokens can aid visual pre-training to form exceptional long-term modeling capabilities. Among these, Titok tokenizer [86] focuses on encoding images with a reduced number of tokens, while the 4M [43] specializes in aligning and jointly training across multiple modalities. MAGVIT [84, 85] uses a super-large codebook based on a look-up-free quantization method.
3 Method
3.1 Background
Video continuation.
We define the video continuation task as generating the future frames $\{\tilde{I}_{t+1},\tilde{I}_{t+2},\tilde{I}_{t+3},\ ...\}$ given a sequence of past $t$ frames $\{I_{i}\in\mathbb{R}^{h\times w\times 3}|_{i=1,...,t}\}$. The inputs of our video continuation model are the three consecutive frames $I_{t-2}$, $I_{t-1}$, and $I_{t}$. Our auto-regressive paradigm allows us to generate an arbitrary number of frames iteratively. Our ARCON approach can generate creative minute-level videos, based on the model design paradigm of generating video frames in an auto-regressive manner. Our model comprises three decoupling components: (1) an image tokenizer encoding images and semantic maps into discrete tokens, (2) an LVM trained with the next token prediction task to perform the video continuation task in an auto-regressive manner, and (3) an image decoder that can decode discrete tokens to images.
Image tokenizer.
We choose the MAGVIT-v2 tokenizer [85] to encode the RGB images and semantic maps. We hope to keep the encoder of the tokenizer unchanged because such modifications would require offline reprocessing of the entire training set and would also affect the subsequent training of the generative model. We directly utilize the open-source weights provided by Open-MAGVIT2 [41]. By leveraging a lookup-free quantization approach, this tokenizer achieves impressive image reconstruction results with an extremely large vocabulary. As a result, this tokenizer will tokenize a $112\times 112\times 3$ image into $14\times 14$ discrete tokens, with a vocabulary size of $2^{18}$. We find that this compression ratio is already quite extreme, and it is difficult to reduce the image quality loss caused by encoding and decoding under this bottleneck. We will discuss in the subsequent sections a more direct integration of the input frame texture within the decoder part.
LVM.
Based on MAGVIT-v2’s token factorization technique [85], instead of predicting using a codebook of size $2^{18}$, we can predict using two concatenated codebooks, each of size $2^{9}$. The final result is that each $112\times 112\times 3$ frame will be converted into $784$ 1D tokens with a vocabulary size of $2^{9}$. We adopt the LLaMA [59] architecture which is a popular open-source model. We train the model from scratch with a context length of $16,384$ tokens, which can fit no more than $20$ images under MAGVIT-v2 [85] tokenizer. We package the data into a question-and-answer format and use a system prompt to specify whether the task is to continue with pure RGB tokens or to interleave semantic/RGB tokens. Once frames can be represented as token sequences, we can train the model to minimize the cross-entropy loss for predicting the next token.
3.2 Interleaving generation
The benefit of the auto-regressive framework is that multimodal data from a variety of data structures (text, images, videos, audio, etc.) can all be converted into discrete tokens so that language models can understand the information from the various modalities based on different codebooks. Many works have attempted to utilize this approach for translation between multimodal data [2, 32], or for image generation using auxiliary modalities (depth maps, segmentation maps, sketches, etc.) as control signals [82]. Auxiliary modalities can carry much structural information of an image at an information density much lower than that of an RGB image. We believe that generating sequences of auxiliary modal tokens along with sequences of RGB tokens helps the LVM simplify the video continuation task. By incorporating the task of continuing the auxiliary modal sequences, we essentially break down the video continuation task into two sub-tasks: continuing the auxiliary modality and translating between modalities. This decomposition allows the model to capture structural details with fewer tokens, while also enabling the generation of new objects and scenes that were not present in the input sequence. This approach reduces the computational cost for the model and enhances both the temporal consistency and creative flexibility of the video continuation. Similar ideas have also demonstrated effectiveness in recent speech generation work, Step-Audio [28].
The structure of our ARCON model is shown in Fig. 2. First, we extract the semantic segmentation maps of the video data using a pre-trained Uniformer [33]. Then we utilize an image tokenizer that can encode images from RGB pixel space to discrete token space. We use this tokenizer to extract RGB tokens from video frames, as well as semantic tokens from the corresponding auxiliary modalities. Finally, we employ an LVM based on LLaMA [59], in its training and inference phases, we interleave the semantic tokens and the RGB tokens, so that the model generates semantic tokens and then RGB tokens at each timestamp to achieve the video continuation task.
Figure 3: Flow-based feature warping in decoder. During the decoding of generated tokens, some auxiliary features can be transferred from reference tokens using a flow-based warping. The input to the right-side decoder is additionally concatenated with a warped feature. We can opt for higher-resolution reference frames to provide larger feature maps, and the feature maps of the generated frames are aligned with them through bilinear resizing.
3.3 Token decoder
Representing high-definition images with very few discrete tokens inevitably leads to quality loss [85, 52]. To mitigate this, previous work [25, 37] has demonstrated that training a large decoder, such as a video diffusion decoder, can remember various texture details. In our work, we investigate the use of a simple method inspired by reference-based super resolution [90, 39] to borrow texture information from high-resolution input frames, thereby enhancing video quality. We train the flow-based feature warping method on BDD100K using open-MAGVIT-v2 scheme [41]. We keep the encoder frozen and fine-tune the MAGVIT-v2 decoder. This fine-tuning approach enhances the quality of reconstructed videos in an offline manner.
Fig. 3 is a schematic diagram illustrating our decoder. A simple flow model predicts a flow f that aligns between two feature maps using several layers of $3\times 3$ convolution layers. We use the reference tokens from high-resolution input frames with the shape of $224\times 224\times 3$ during inference.
4 Experiments
This section conducts experiments on our discrete token-based auto-regressive model and verifies the suitability of the proposed image tokenizer and LVM for long-term video continuation tasks. Quantitative and qualitative results demonstrate that semantic tokens can help generate more creative and temporally consistent long videos.
4.1 Experimental setup
Modality Coverage rate of training data
50%       95% 99%
RGB codebook rate 23%       82% 94%
Semantic codebook rate 1%       30% 54%
Table 1: Usage ratio of codebooks across different modalities. The codebook size is $2^{18}$. Compared to RGB tokens, semantic tokens use less codebook space to cover more training data.
Datasets.
We choose BDD100K [83] as our model’s training data, a large-scale dataset of autonomous driving scenarios. It contains $100,000$ videos, most of which are at $30$ FPS, with video duration generally in the $30$ to $40$ seconds range. In the training stage, We only use its official training set, which contains $70,000$ videos totaling about $1,100$ hours. We believe that it has sufficient data diversity to allow our video continuation model to obtain strong generalization within the autonomous driving domain, and it has an appropriate average video duration to explore how the model exhibits creativity in long-term video generation. We sample all videos at $3$ Hz and center crop all frames to $112\times 112\times 3$, then encode them to $392$ tokens per image using our image tokenizer, so the total training data we use contains about $3$B tokens. We evaluate the video continuation task at $0.6$ FPS in the inference stage. We select $100$ random video clips from the BDD100K [83] test set and all $150$ videos from the nuScenes [7] validation set for testing.
We utilize Kinetics-700 (K700)  [8, 9] dataset for ablation experiments on Tokenizer. K700 is a large-scale video dataset with extensive action category annotations, providing highly diverse and high-quality videos. We evaluated the effectiveness of the flow-based feature warping method on the validation set of K700.
Training procedure.
Each training sample used by our auto-regressive model contains tokens from $18$ consecutive frames. Half of the training samples consist solely of RGB tokens, while the remaining samples are interleaved with both semantic tokens and RGB tokens. The global batch size is set as $64$. We use a linear annealing strategy to reduce the learning rate from $10^{-5}$ to $2\times 10^{-6}$. The $7$B probing model is trained for $20K$ iterations on $256$ A800 GPUs, which takes about $12$ hours. The total number of image tokens in the training experience is approximately $14$B. For the 20B model, we triple the number of layers from the 7B model, and we find that the larger model quickly learns the correct number of output tokens and a stable format. We train the $20$B model for $50K$ iterations with other settings unchanged which takes about $90$ hours. The MAGVIT-v2 is fine-tuned for $200$ epochs on the BDD100K training set, which takes about $6$ hours.
4.2 Probing Experiments
We report our findings from some probing experiments with the ARCON model of 7B.
Figure 4: Optical flow vector magnitude decay. We generate 150 15-frame clips on the nuScenes validation set and compute the optical flow mean magnitude with a pre-trained RAFT model [57] between adjacent frames. A lower value indicates less motion, i.e., a more stationary video. Results demonstrate that continuing semantic maps helps mitigate degradation in video generation.
Figure 5: Consistency between generated semantic maps and RGB images. We use the same Uniformer [33] model used in the pipeline to perform semantic segmentation on the frame sequence generated by our ARCON model. We confirm there is a high degree of correspondence when the two modalities are generated alternately. This indicates that this approach implicitly decomposes the video sequencing task into a semantic sequencing task and a semantic map to the RGB translation task. Thanks to the auto-regressive paradigm, this translation task is video-consistent.
Figure 6: Video continuation samples. The first example demonstrates that our model can make autonomous decisions about the ego car’s driving action by the driving action of the car in front of it and that the car in front of it will still appear in the picture after making a turn. The second example demonstrates that there is a small probability that the model will choose to turn around and leave when the car in front slows down and stops. The third example shows that the brake lights come on at the right time when the car in front is slowing down. The last example shows that our model can generate completely new scenes, even learning the physical knowledge that the reflection on the hood should be consistent with the in-frame picture.
Codebook usage across different modalities.
As shown in Tab. 1, semantic tokens require a smaller vocabulary space. Upon analyzing the token utilization across all BDD100K [83] training data, it is observed that semantic tokens occupy merely 1% of the codebook space to encompass 50% of the training data, in contrast to RGB tokens which utilize 23% of the codebook space. Notably, half of the codebook space suffices to cover 99% of semantic videos, whereas RGB videos necessitate 94% of the codebook space. These metrics demonstrate the semantic tokens at a higher level of abstraction.
Semantic token first.
We believe alternatively generating semantic tokens and RGB tokens makes the auto-regressive model pay more attention to the structure of the videos. Placing semantic tokens before RGB tokens essentially breaks down the video continuation task into continuation and translation subtasks. It allows the model to capture structural details explicitly. We posit that incorporating semantic tokens helps auto-regressive models preserve video structural information. We find incorporating a semantic token generation step before image token generation within an auto-regressive model can improve long-term generation capabilities through mitigating the degeneration phenomenon. To quantify this phenomenon, here we propose to use the average optical flow vector magnitude between neighboring frames to quantify the overall motion of the video. Results are shown in Fig. 4.
Consistency between two modalities.
We find that interleaving generated semantic maps and RGB images have a high degree of consistency, results shown in Fig. 5. We use the same semantic segmentation model to re-extract the semantic segmentation maps for the generated images and find that the accuracy of the semantic segmentation maps generated by the auto-regressive model can reach 77.4% on average on nuScenes validation set.
Figure 7: The influence of semantic tokens. We choose a challenging nighttime scenario to highlight the importance of semantic segmentation, as it is less affected by the lighting conditions in the video. Semantic tokens assist the model in regenerating the black vehicle ahead after the left turn.
Figure 8: Generation across multiple timelines. The leftmost arrows indicate the driving directions of the green car ahead and the ego car, respectively.
Methods FID $\downarrow$ FVD $\downarrow$
nuScenes fine-tuning
DrivingDiffusion [34] 15.6 335
DriveDreamer [67] 52.6 452
WoVoGen [38] 27.6 417.7
SubjectDrive [29] 16.0 124
Panacea [73] 17.0 139.0
Drive-WM [70] 15.8 122.7
DriveDreamer-2 [87] 18.4 74.9
Vista [16] 6.9 89.4
w/o nuScenes fine-tuning
DriveGAN [31] 73.4 502.3
GenAD (OpenDV-2K) [79] 15.4 184.0
DrivingWorld [27] 16.4 174.4
LVM [2] $\dagger$ 71.5 162.2
StreamingT2V [24] $\dagger$ 28.8 131.0
CogVideoX [81] $\dagger$ 64.9 79.8
Ours 23.3 57.6
Table 2: Comparison with other methods on nuScenes validation set. $\dagger$ denotes general video generation algorithms. We achieve the best FVD results on the nuScenes dataset without the need for fine-tuning [7].
4.3 Quantitative results
We evaluate our ARCON model’s capability on video continuation and compare with related video generation methods. As exhibited in Tab. 2, the quantitative results show that even without training or fine-tuning on the nuScenes dataset [7], our model still outperforms other baselines on Fréchet Video Distance (FVD) scores [60], which proves that our model is not only capable of generating high-quality videos but also has strong generalization ability.
4.4 Qualitative results
Unless specified, all of our visualization results are conditioned on the first $3$ frames, and sampled from the BDD100K dataset at $0.6$Hz.
Highly creative generation.
Our model possesses the capability to continuously generate scenes from the perspective of a moving vehicle. Notably, the model exhibits an autonomous learning process, acquiring fundamental traffic knowledge. For instance, it rarely tries to turn left in the right turn lane, and demonstrates a tendency to decelerate or maneuver around vehicles positioned in front of it. Furthermore, the model consistently generates coherent traffic lights, lane markings, signs, and crosswalks within urban contexts. The outcomes of these capabilities are visually represented in Fig. 6.
Multiple timelines.
Our model exhibits the capability to generate a multitude of diverse future outcomes. Given the same 3-frame input, we randomly produce $50$ distinct future videos. In the specific scenario where the vehicle ahead intends to turn right, our model tends to follow the lead vehicle and turn right as well. However, there exists a 34% probability that the ego vehicle will proceed straight ahead. Additionally, in 6% of the generated videos, the vehicle ahead temporarily opts to go straight. These outcomes are visually presented in Fig. 8.
Auto-regressive iteration.
Due to the alternating generation of semantic and RGB video frame sequences, our ARCON model is capable of auto-regressively generating minute-level videos. Results are shown in Fig. 1.
As shown in Fig. 7, the qualitative results demonstrate that utilizing semantic tokens can help the video continuation model to generate results with stronger video consistency, while making the results not easily degrade to copies of the last frame due to the semantic tokens prediction being more sensitive to small motion in the video frame sequences. As exhibited in Tab. 2, the quantitative results show that even without training or fine-tuning on the nuScenes dataset [7], our model still outperforms other baselines on Fréchet Video Distance (FVD) scores [60], which proves that our model is not only capable of generating high-quality videos but also has strong generalization ability. In Fig. 9, we compare our ARCON model with general video generation models.
Figure 9: Video continuation. As an optical flow-based method, DMVFN [26] struggles to generate objects or scenes that do not appear in historical frames. Meanwhile, CogVideoX [81] tends to produce motion blur in areas with significant motion. In contrast, our ARCON model generates coherent new scenes while better preserving the structural information of objects.
4.5 Ablation study
As shown in Tab. 3, we investigate the efficacy of our simple flow-based feature warping method in the decoding process. We report the FVD metric on the in-distribution BDD100K [83] and out-of-distribution K700 [9] datasets. Cross-frame feature transfer relies on the similarity between frames, with lower similarity increasing the difficulty of the transfer. We report the FVD metric at various frame rates, which result in differing inter-frame similarities. FVD metrics are evaluated on $400$ samples in validation split with $16$-frame clips at the resolution of $224\times 224$. The method has demonstrated benefits in both in-distribution scenarios and out-of-distribution scenarios.
Setting FVD $\downarrow$ on BDD100K
30FPS 10FPS 3FPS
MAGVIT-v2 506.0 261.2 90.7
+ Feature warping 249.5 139.2 62.5
Setting FVD $\downarrow$ on K700
30FPS 10FPS 3FPS
MAGVIT-v2 538.4 233.4 63.8
+ Feature warping 308.3 146.3 93.4
Table 3: FVD metrics on BDD100K [83] dataset and K700 [9] dataset at different frame-rates. Under higher frame rate tests, the magnitude of FVD will increase, which, although not intuitive, can be corroborated by the metrics of related works [61, 80].
We further analyze two factors that are most likely to affect the creativity of the video continuation model, including the additional modalities utilized, and the choice of temperature during inference in Tab. 4.
Setting BDD100K nuScenes
FID $\downarrow$ FVD $\downarrow$ FID $\downarrow$ FVD $\downarrow$
semantic segmentation
w/o sem 35.4 91.6 30.4 84.2
baseline 29.4 73.6 23.3 57.6
inference temperature $t$
$t=0.2$ 38.6 94.9 32.2 106.1
$t=0.5$ 32.4 75.2 27.3 70.4
$\mathbf{t=0.7}$ 29.4 73.6 23.3 57.6
$t=1.0$ 44.2 150.3 34.4 129.1
Table 4: The impact of interleaved generation and inference temperature on quantitative metrics.
5 Conclusion
We develop an ARCON scheme that interleaves the generation of RGB tokens and semantic tokens. We adhere to a generative paradigm that separates structure from texture, demonstrating the effect of semantic tokens on the generation of RGB tokens. Our approach can be improved at several aspects. At present, the representation of a single image necessitates hundreds of tokens, combined with the extensive parameter count of LVM, leading to notably slow inference speeds. Novel encoding strategies [88, 15] can offer enhancements to the efficiency of this approach. Additionally, our texture transfer method lacks effective guidance for long-term generation tasks. Future research could investigate the integration of flow-based techniques with generative methods at the decoding stage [42]. Furthermore, exploration into the physical consistency of generative models is pivotal for autonomous driving applications.